# What Has Been Lost with Synthetic Evaluation?

### v2

@@ Abstract
Large language models (LLMs) are increasingly used for data generation. However, creating evaluation benchmarks raises the bar for this emerging paradigm. Benchmarks must target specific phenomena, penalize exploiting shortcuts, and be challenging. Through two case studies, we investigate whether LLMs can meet these demands by generating reasoning-over-text benchmarks and comparing them to those created through careful crowdsourcing. Specifically, we evaluate both the validity and difficulty of LLM-generated versions of two high-quality reading comprehension datasets: CondaQA, which evaluates reasoning about negation, and DROP, which targets reasoning about quantities. We find that prompting LLMs can produce variants of these datasets that are often valid according to the annotation guidelines, at a fraction of the cost of the original crowdsourcing effort. However, we show that they are less challenging for LLMs than their human-authored counterparts. This finding sheds light on what may have been lost by generating evaluation data with LLMs, and calls for critically reassessing the immediate use of this increasingly prevalent approach to benchmark creation.
@@ 1 Introduction
The development of benchmarks that test frontier models in reasoning over text has been at the heart of NLP research Rogers et al. (2023). Such benchmarks aim to cover a broad range of input variations and related skills, and prevent models from performing well by exploiting data shortcuts Bowman and Dahl (2021). They are created to be challenging, so that even the most capable emerging models cannot solve them immediately and they remain a meaningful yardstick for measuring progress. Yet these goals are often hindered by crowdsourcing annotation practices that favor speed to maximize hourly pay, such as crowdworkers reusing the same pattern(s) to create dataset examples. An alternative approach, prompting LLMs, does not suffer from the same misaligned incentives. LLMs can create content they struggle to reason about (West et al., 2024, the generative AI paradox). However, the challenge now lies in getting LLMs to follow annotation instructions. The silver lining is that it is possible to review annotations and revise guidelines without delay because LLMs produce annotations instantly. Moreover, since their annotations come at little or no cost, one can iterate as many times as they are willing to review-and-revise.
This makes it tempting to substitute LLMs for human annotators, as is increasingly common (§3). However, before adopting this approach to creating new challenging benchmarks, we must ask: When created under the same annotation criteria, how do LLM-generated benchmarks compare to their challenging human-authored counterparts? If the data is not comparable in quality or as effective for assessing model limitations, this should give us pause in using LLMs as annotators in this context.
We address this question through the lens of reading comprehension, which is a common format in NLP for probing reasoning over text capabilities of frontier models Rogers et al. (2023). Specifically, we study compositional reasoning over numbers as in DROP Dua et al. (2019); Gardner et al. (2020) and reasoning about the implication of negation as in CondaQA Ravichander et al. (2022). These reasoning types continue to challenge LLMs, as shown in Table 1.
@@
These findings suggest that the core challenge is not validity, but difficulty. However, optimizing prompts for difficulty would require substantial human-in-the-loop effort because LLMs cannot yet reliably validate and answer their own generations. Moreover, this would effectively turn evaluation-data generation into a form of adversarial dataset construction whose “naturalness” has been questioned Bowman and Dahl (2021). Therefore, how to generate challenging instances without relying on excessive human effort and producing artificially-sounding text remains an open question. In light of this, using LLMs as annotators for any new benchmark does not seem advisable as of yet, especially if the validity of the generated data is the only measure of dataset quality.
@@ 2 Background
See Table E (Appendix) for examples of human-authored questions in both datasets. Questions in these datasets are posed over Wikipedia passages. In DROP, questions must “require complex linguistic understanding and discrete reasoning” Dua et al. (2019). An extra condition is that the BiDAF model Seo et al. (2017) cannot answer a question. We do not adversarially generate, as this may raise other issues Bowman and Dahl (2021). In CondaQA, questions should be “targeted towards the implications of the negated statement” Ravichander et al. (2022), i.e, they should be about (1) the negated statement, rather than other information in the passage, and (2) about an implication of the negated statement.
@@
Models can solve benchmarks by exploiting data shortcuts Gururangan et al. (2018). Thus, they are better evaluated based on consistency — their ability to solve an entire set of slightly different examples with different ground truth, known as contrast sets Gardner et al. (2020) or counterfactual data Kaushik et al. (2020). Contrastive edits are valid if minimal changes give a different ground truth, while maintaining fluency. However, to assess reasoning capabilities, edits targeting specific changes help evaluate whether models have the necessary skills for that type of reasoning. For example, an edit that requires changing only what is negated tests models’ understanding of the negation scope:
@@
Only with this ability can models consistently reason and answer questions as “Was Parsons able to use his guitar after the fire?” in both contexts. We introduce other edit types below.
@@
Gardner et al. (2020) report three DROP edit types. Compositional edits add a reasoning step to questions (Table E; Appendix). Semantic edits invert question meanings, such as changing “shortest” to “longest”. Temporal edits modify event order by changing their order in the questions or the passage (Table 7; Appendix), or by changing the dates associated with these events in the passage. Unfortunately, the DROP contrast set does not explicitly mark the edit type. [^3] https://github.com/allenai/contrast-sets/blob/main/DROP/drop_contrast_sets_test.json Thus, one author of this paper has manually determined compositional and temporal edits that change event order in passages, but could not reliably detect semantic edits. [^4] We choose passage-level temporal edits because they with compositional edits cover at least one type of both question-level and passage-level DROP edits. We generate compositional edits using pairs of passages and the generated questions. Because the contrast set contained too few temporal edits, we generate temporal edits using the original DROP dev passages that include dates.
@@ 3 Related Work
Figure 1: The percentage of valid questions in a sample of 16 for intermediate evaluations, and 100 or 114 for final evaluations. Questions are generated with gpt-4-turbo-2024-04-09 and assessed by one author of this paper.
@@
A line of work has aimed to automatize the creation of contrastive edits to improve their scale and diversity (Li et al., 2020; Ross et al., 2021; Shaikh et al., 2022; Ross et al., 2022; Dixit et al., 2022; Chen et al., 2023, among others); reviewed in Stepin et al. (2021); Nguyen et al. (2024); Wang et al. (2024). While these approaches accept any minor change that alters the answer, we focus on editing that is done along specific dimensions such as changing what is negated or adding a set of calculations. Ross et al. (2022) also steer edits toward targeted attributes, but their approach requires dataset creators to adjust semantic representation-based control codes, whereas we achieve targeted edits through direct natural language instructions.
@@
A new trend is to generate data with LLMs Long et al. (2024), spanning NLP tasks such as text classification Møller et al. (2024); Patwa et al. (2024), QA Ye et al. (2022); Harsha et al. (2025), NLI Liu et al. (2022); Hosseini et al. (2024), as well as instruction tuning and alignment Wang et al. (2023); Bai et al. (2022). Our work differs from this line of work in three key ways. First, most prior efforts focus on generating training, not evaluation, data. Imperfect synthetic data may improve training Amin et al. (2025), but is unacceptable for benchmarking. Second, much of the existing work targets the easier task of generating only labels or answers, whereas we generate content: questions and edits. Third, our goal is to critically and systematically analyze generated data. While some studies generate evaluation datasets or replicate existing crowdsourcing pipelines Gilardi et al. (2023); Qi et al. (2025); Roemmele and Gordon (2024); Wu et al. (2025), they do not conduct such analysis. Related are Das et al. (2024), who study artifacts in LLM-generated data but focus on different data types, and Maheshwari et al. (2024), who analyze generated labels rather than new content.
@@ 4 Generating DROP and CondaQA
We give a Wikipedia passage from the original dataset to gpt-4-turbo-2024-04-09 and refine prompt instruction iteratively, with one of the authors manually validating 16 generations for each promising candidate prompt. [^5] This was the state-of-the-art LLM at the time. We consider an annotation valid if it adheres to the original annotation guidelines; see validity criteria in Table 11 (Appendix). If validity is low, we adjust the prompt until we reach a promising version, which we then re-evaluate on the 16 examples. Upon achieving either 100% validity, or if further prompt engineering suggests no prospects of progress, the same author evaluated the most valid intermediate prompt on a larger sample of 100+ instances. Each iteration in Figures 12 represents our prompt after exploring several minor variations around a central prompting idea in that iteration. We also show these results in Tables 12 and 13; Appendix. In total, we evaluate 421 paragraph-question pairs and 926 edits across both datasets in this section.
@@ 4.1 Prompting for Question Generation
Figure 2: The percentage of valid edits in a sample of 16 for intermediate evaluations, and 100 or 114 for final evaluations. Edits are generated with gpt-4-turbo-2024-04-09 and assessed by one author of this paper.
@@
For CondaQA, we first introduce filters for (1)–(3) above, but we found including a filter for (4) was crucial (see Iter 4 vs. 5 in Fig. 1). After this iteration, we are unable to further improve the validity of generated CondaQA questions, and ultimately, their validity in a sample of 114 passage-question pairs is 72.8%.
For DROP, we refine the filtering in a few iterations, each of which further improves the validity. The first filter checks that a question requires complex reasoning and looking over more than one part of the passage to answer it (Iter 4). The next filter checks that the question does not combine multiple questions into one and that it has a valid answer type. Before introducing the final filter, we enrich the prompt with an example of 9 types of discrete reasoning questions and request generating a question for each discrete reasoning type, then ranking them according to how well they adhere to the other guidelines (see Iter 5). [^8] Types of discrete reasoning Dua et al. (2019): subtraction, comparison, selection, addition, count and sort, coreference resolution, other arithmetic, set of spans, and other. Given this improvement, we try a similar addition for CondaQA (Iter 6), but it drastically reduces validity. [^9] Dimensions of commonsense reasoning Ravichander et al. (2022): precondition, social norms, psychology, cause and effect, mutual exclusivity, synecdoche. This demonstrates that a prompt engineering strategy crucial for producing one reasoning benchmark may not be universal. The final filter checks that questions do not require any outside knowledge (e.g., about a sport or player stats) to be answered. With this final filter, we reach 100% validity on the small sample which leads to 88.7% valid generated DROP questions in a larger sample of 100 passage-question pairs. This performance drop from a small- to large-scale evaluation pinpoints a limitation of this protocol, which could also happen in traditional human evaluation.
@@ 4.2 Prompting for Editing
In most cases, the model needs to be primed to focus on relevant information and guided through a multi-step generation process. For instance, in affirmative edits, the initial prompt (Iter 3) involves removing the negation cue and then polishing the text. A more effective alternative (Iter 5) is to explicitly ask what is being negated, de-negate the negation while keeping the original negated content in focus, and then integrate the revised content into the passage. We often introduce overgenerating-and-filtering with other refinements, which makes its role in editing less immediately apparent. However, it is crucial (except for paraphrasing), as evident by using it in all final prompts.
@@ 5 Whose Annotations are Preferred?
We next study whether NLP researchers prefer human or generated questions and edits based on how closely they follow the original guidelines for producing questions and edits with desired properties. Before we proceeded, we eliminated the possibility that generations are paraphrased instances from the human-authored counterparts; see Appendix B.
@@
Screenshots of this study are shown in Figures 411 for CondaQA and Figures 1825 for DROP; in the Appendix. Annotators were asked to imagine themselves as dataset creators who were designing a benchmark to probe LLMs’ capabilities to reason and were provided background about the format of the datasets (without mentioning their name). They were then asked to choose edits and questions that better meet the dataset specification. They were also asked if the question or edit was invalid. For CondaQA, annotators are shown one question pair and three edit pairs corresponding to the three edit types. For DROP, annotators are shown one question pair and one compositional-edit pair. [^12] We excluded temporal edits because we could identify only $\approx$20 human-authored ones. Since the compositional edits in §4 were made on generated questions, we created compositional edits for human-authored questions here to enable a valid comparison. The order of human and generated annotations for both studies is randomized. Other details are in Appendix §C.
@@
Table 2: The validity assessments for questions and edits from the human preferences study.
@@ Results.
Tables 2 and 3 show the results. The caption of Table 3 explains how the results are structured. We observe that annotators assess both generated and human edits and questions as valid at a much lower rate than one author of this paper, with the exception of DROP compositional edits. They also rate the original CondaQA paraphrase/scope edits and questions, along with the DROP questions, as less valid than the generated ones. Moreover, annotators are often shown to prefer model-generated questions and edits to human-written questions and edits from the original dataset.
This likely stems from the fact that the generated edits tend to be simpler and, due to the way the prompts were designed, follow the guidelines for each type of edit more closely. In contrast, original annotators often interpret the question and edit definitions in slightly different ways and sometimes introduce extra flair to the annotations. While these human annotations are, in our opinion, still valid, they do not always strictly follow the edit guidelines. This is aligned with observations from prior work Ruggeri et al. (2024). However, when juxtaposed with the generated edits and evaluated by NLP researchers based solely on the edit definitions, the generated annotations’ stricter compliance with the guidelines points out some of the smaller flaws in the human edits that would not necessarily be evident if one only saw those edits in isolation.

### v3

@@ Abstract
Large language models (LLMs) are increasingly used for data generation. However, creating evaluation benchmarks raises the bar for this emerging paradigm. Benchmarks must target specific capabilities, penalize exploiting shortcuts, and be challenging. Through two case studies, we investigate whether LLMs can meet these demands by generating reasoning-over-text benchmarks and comparing them to those created through careful crowdsourcing. Specifically, we evaluate both the validity and difficulty of LLM-generated versions of two high-quality reading comprehension datasets: CondaQA, which evaluates reasoning about negation, and DROP, which targets reasoning about quantities. We find that prompting LLMs can produce variants of these datasets that are often valid according to the annotation guidelines, at a fraction of the cost of the original crowdsourcing effort. However, we show that they are less challenging for LLMs than their human-authored counterparts, do not preserve the same model rankings, and that these differences may be imperceptible to researchers inspecting the data. This finding sheds light on what may have been lost by generating evaluation data with LLMs, and calls for critically reassessing the immediate use of this increasingly prevalent approach to benchmark creation.
@@ 1 Introduction
The development of benchmarks that test frontier models in reasoning over text has been at the heart of NLP research Rogers et al. (2023). Such benchmarks aim to cover a broad range of input variations and related skills, and prevent models from performing well by exploiting data shortcuts Bowman and Dahl (2021). They are created to be challenging, so that even the most capable emerging models cannot solve them immediately and they remain a meaningful yardstick for measuring progress. Yet these goals are often hindered by crowdsourcing annotation practices that favor speed to maximize hourly pay, such as crowdworkers reusing the same pattern(s) to create dataset examples. An alternative approach, prompting LLMs, does not suffer from the same misaligned incentives. LLMs are capable of creating content they will struggle to reason about (West et al., 2024, the generative AI paradox). However, a key challenge lies in getting LLMs to follow annotation instructions. The silver lining is that it is possible to review annotations and revise guidelines without delay because LLMs produce annotations instantly. Moreover, since their annotations come at little or no cost, one can iterate as many times as they are willing to review-and-revise.
This makes it tempting to substitute LLMs for human annotators, as is becoming increasingly common (§3). However, before adopting this approach to creating new challenging benchmarks, we must ask: When created under the same annotation criteria, how do LLM-generated benchmarks compare to their challenging human-authored counterparts? If the data is not comparable in quality or as effective for assessing model limitations, this should give us pause in using LLMs as annotators in this context.
We address this question through the lens of reading comprehension, which is a common format in NLP for probing reasoning over text capabilities of frontier models Rogers et al. (2023). Specifically, we study compositional reasoning over numbers as in DROP Dua et al. (2019); Gardner et al. (2020) and reasoning about the implications of negated statements as in CondaQA Ravichander et al. (2022). These reasoning types continue to challenge LLMs, as shown in Table 1.
@@
These findings suggest that the core challenge is not validity, but difficulty. However, optimizing prompts for difficulty would require substantial human-in-the-loop effort because LLMs cannot yet reliably validate and answer their own generations. Moreover, this would effectively turn evaluation-data generation into a form of adversarial dataset construction whose “naturalness” has been questioned Bowman and Dahl (2021). Therefore, how to generate challenging instances without relying on excessive human effort and producing artificially-sounding text remains an open question. In light of this, using LLMs as annotators for any new benchmark does not seem advisable as of yet, especially if the validity of the generated data is the only measure of dataset quality. [^1] Our code is available at Github.
@@ 2 Background
See Table 8 in the Appendix for examples of human-authored questions in both datasets. Questions in these datasets are posed over Wikipedia passages. In DROP, questions must “require complex linguistic understanding and discrete reasoning” Dua et al. (2019). An extra condition is that the BiDAF model Seo et al. (2017) cannot answer a question. We do not adversarially generate, as this may raise other issues Bowman and Dahl (2021). In CondaQA, questions should be “targeted towards the implications of the negated statement” Ravichander et al. (2022), i.e, they should be about (1) the negated statement, rather than other information in the passage, and (2) about an implication of the negated statement.
@@
Models can solve benchmarks by exploiting data shortcuts Gururangan et al. (2018). Thus, they are better evaluated based on consistency — their ability to solve an entire set of slightly different examples with different ground truth, known as contrast sets Gardner et al. (2020) or counterfactual data Kaushik et al. (2020). Contrastive edits are valid if minimal changes give a different ground truth, while maintaining fluency. However, edits targeting specific changes help evaluate whether models have the necessary skills for a specific type of reasoning. For example, an edit that requires changing only what is negated tests models’ understanding of the negation scope:
@@
Only with this skill can models consistently reason about negation and answer questions as “Was Parsons able to use his guitar after the fire?” in both contexts. We introduce other edit types below.
@@
Gardner et al. (2020) report three DROP edit types. Compositional edits add a reasoning step to questions (Table 8; Appendix). Semantic edits invert question meanings, such as changing “shortest” to “longest”. Temporal edits modify event order by changing their order in the questions or the passage (Table 6; Appendix), or by changing the dates associated with these events in the passage. Unfortunately, the DROP contrast set does not explicitly mark the edit type. [^4] https://github.com/allenai/contrast-sets/blob/main/DROP/drop_contrast_sets_test.json Thus, one author of this paper has manually determined compositional and temporal edits that change event order in passages, but could not reliably detect semantic edits. [^5] We choose passage-level temporal edits because they, with compositional edits, cover at least one type of both question-level and passage-level DROP edits. We generate compositional edits using pairs of passages and the generated questions. Because the contrast set contained too few temporal edits, we generate temporal edits using the original DROP dev passages that include dates.
@@ 3 Related Work
Figure 1: The percentage of valid questions in a sample of 16 for intermediate evaluations, and 100 or 114 for final evaluations. Questions are generated with gpt-4-turbo-2024-04-09 and assessed by one author of this paper. The full prompt code and templates for each iteration can be found in our Github repository.
@@
A line of work has aimed to automate the creation of contrastive edits to improve their scale and diversity (Li et al., 2020; Ross et al., 2021; Shaikh et al., 2022; Ross et al., 2022; Dixit et al., 2022; Chen et al., 2023, among others); reviewed in Stepin et al. (2021); Nguyen et al. (2024); Wang et al. (2024). While these approaches accept any minor change that alters the answer, we focus on editing that is done along specific dimensions such as changing what is negated or adding a set of calculations. Ross et al. (2022) also steer edits toward targeted attributes, but their approach requires dataset creators to adjust semantic representation-based control codes, whereas we achieve targeted edits through direct natural language instructions.
@@
Generating data with LLMs is increasingly common Long et al. (2024), from NLP tasks such as classification Møller et al. (2024); Patwa et al. (2024), QA Ye et al. (2022); Harsha et al. (2025), NLI Liu et al. (2022); Hosseini et al. (2024), to instruction tuning and alignment Wang et al. (2023); Bai et al. (2022). Our work differs from this line of work in three key ways. First, most prior efforts focus on generating training, not evaluation, data. Imperfect synthetic data may improve training Amin et al. (2025), but is unacceptable for benchmarking. Second, much of the existing work targets the easier task of generating only labels or answers Maheshwari et al. (2024), whereas we generate content: questions and edits. This is more challenging to execute due to more complex prompt engineering and validity review, as well as costly answer annotation. Third, our goal is to critically and systematically analyze generated data. While some studies generate evaluation datasets or replicate existing crowdsourcing pipelines Gilardi et al. (2023); Qi et al. (2025); Roemmele and Gordon (2024); Wu et al. (2025), they do not conduct such analysis. Das et al. (2024) compares LLM-generated data to human-authored counterparts, but focuses on training data and the downstream effects of its use. Our results both corroborate and go beyond Maheshwari et al. (2024)’s. Specifically, we show that synthetic data is not always a good predictor of relative performance, and that bias does not explain why synthetic benchmarks are easier.
@@ 4 Generating DROP and CondaQA
We give a Wikipedia passage from the original dataset to gpt-4-turbo-2024-04-09 (the state-of-the-art LLM at the time) and refine prompt instruction iteratively, with one of the authors manually validating 16 generations for each promising candidate prompt. We consider an annotation valid if it adheres to the original annotation guidelines; see validity criteria in Table 11 in the Appendix. If validity is below 85%, we adjust the prompt and experiment using a chat interface until we reach a promising version, which we then re-evaluate on the 16 examples. Upon achieving either 100% validity, or if further prompt engineering suggests no prospects of progress, the same author evaluated the most valid intermediate prompt on a larger sample of 100+ instances. Each iteration in Figures 12 represents our prompt after exploring several minor variations around a central prompting idea in that iteration. We also show these results in Tables 12 and 13; Appendix. In total, we evaluate 421 paragraph-question pairs and 926 edits across both datasets in this section.
@@ 4.1 Prompting for Question Generation
Figure 2: The percentage of valid edits in a sample of 16 for intermediate evaluations, and 100 or 114 for final evaluations. Edits are generated with gpt-4-turbo-2024-04-09 and assessed by one author of this paper. The full prompt code and templates for each iteration can be found in our Github repository.
@@
For CondaQA, we first introduce filters for (1)–(3) above, but we found that including a filter for (4) was crucial (see Iter 4 vs. 5 in Fig. 1). After this iteration, we are unable to further improve the validity of generated CondaQA questions, and ultimately, their validity in a sample of 114 passage-question pairs is 72.8%.
For DROP, we refine the filtering in a few iterations, each of which further improves the validity. The first filter checks that a question requires complex reasoning and looking over more than one part of the passage to answer it (Iter 4). The next filter checks that the question does not combine multiple questions into one and that it has a valid answer type. Before introducing the final filter, we enrich the prompt with an example of 9 types of discrete reasoning questions and request generating a question for each discrete reasoning type, then ranking them according to how well they adhere to the other guidelines (see Iter 5). [^8] Types of discrete reasoning Dua et al. (2019): subtraction, comparison, selection, addition, count and sort, coreference resolution, other arithmetic, set of spans, and other. Given this improvement, we try a similar addition for CondaQA (Iter 6), but it drastically reduces validity. [^9] Dimensions of commonsense reasoning Ravichander et al. (2022): precondition, social norms, psychology, cause and effect, mutual exclusivity, synecdoche. This demonstrates that a prompt engineering strategy crucial for producing one reasoning benchmark may not be universal. The final filter checks that the passage is self-contained in terms of facts required to answer the generated question (e.g., it does not require external knowledge about a sport or player stats). With this final filter, we reach 100% validity on the small sample which leads to 88.7% valid generated DROP questions in a larger sample of 100 passage-question pairs. This performance drop from a small- to large-scale evaluation pinpoints a limitation of this protocol, which could also happen in traditional human evaluation.
@@ 4.2 Prompting for Editing
In most cases, the model needs to be primed to focus on relevant information and guided through a multi-step generation process. For instance, in affirmative edits, the initial prompt (Iter 3) involves removing the negation cue and then polishing the text. A more effective alternative (Iter 5) is to explicitly ask what is being negated, ask the model to de-negate the negation (e.g., change “rarely” to “often”) while keeping the original negated content in focus, and then integrate the revised content into the passage. We often introduce overgenerating-and-filtering with other refinements, which makes its role in editing less immediately apparent. However, it is crucial (except for paraphrasing), as evident by its use in all final prompts.
@@ 5 Whose Annotations are Preferred?
We next study whether NLP researchers prefer human or generated questions and edits based on how closely they follow the original guidelines for producing questions and edits with desired properties. There is a concern about potential dataset contamination when doing such studies with synthetic data, as models may simply memorize and reproduce the human-authored dataset from their training data. To mitigate against the effects of straightforward reproduction, we analyze the generated data to eliminate the possibility that generations are paraphrased instances of their human-authored counterparts; see Appendix B.
@@
Screenshots of this study are shown in Figures 512 for CondaQA and Figures 1926 for DROP (Appendix). [^12] HTML templates for the study are available at our Github. Annotators were asked to imagine themselves as dataset creators who were designing a benchmark to probe LLMs’ capabilities to reason and were provided background about the format of the datasets (without mentioning their names). They were then asked to choose edits and questions that better meet the dataset specification. They were also asked if the question or edit was invalid. For CondaQA, annotators are shown one question pair and three edit pairs corresponding to the three edit types. For DROP, annotators are shown one question pair and one compositional-edit pair. [^13] We excluded temporal edits because we could identify only $\approx$20 human-authored ones. Since the compositional edits in §4 were made on generated questions, we created compositional edits for human-authored questions here to enable a valid comparison. The order of human and generated annotations for both studies is randomized. Other details are in Appendix §C.
@@
Table 2: The validity assessments for questions and edits from the human preferences study. Each row header states the dataset and edit type evaluated. The first column shows the percentage of total edits or questions evaluated which were marked as valid by annotators. All edits and questions evaluated by annotators were also evaluated by an author of this paper. Those results are shown in the second column.
@@ Results.
Tables 23 show the results. We observe that annotators assess both generated and human edits and questions as valid at a much lower rate than one author of this paper, with the exception of DROP compositional edits. They also rate the original CondaQA paraphrase/scope edits and questions, along with the DROP questions, as less valid than the generated ones. Moreover, annotators are often shown to prefer model-generated questions and edits to human-written questions and edits from the original dataset. We include example agreement and disagreement cases between annotators and one author of this paper for generated instances of CondaQA and DROP questions and edits in Tables 1621 in the Appendix.
The annotators’ higher validity and preference for generations likely stems from the fact that the generated edits tend to be simpler and, due to the way the prompts were designed, follow the guidelines for each type of edit more closely. In contrast, original annotators often interpret the question and edit definitions in slightly different ways and sometimes introduce extra flair to the annotations. While these human annotations are, in our opinion, still valid, they do not always strictly follow the edit guidelines. This is aligned with observations from prior work Ruggeri et al. (2024). However, when juxtaposed with the generated edits and evaluated by NLP researchers based solely on the edit definitions, the generated annotations’ stricter compliance with the guidelines highlights small flaws in the human edits that would not necessarily be evident if one only saw those edits in isolation.